# Injecting Bias in Text-To-Image Models via Composite-Trigger Backdoors

### v1

@@ 8.3 Limitation
9 Related Work
9.1 Bias in T2I Models
Several studies [22, 21, 3, 6, 5] have revealed various types of biases in T2I models. Naik et al. [22] found that DALLE-2 [29] and Stable Diffusion [30] exhibit different bias representation ratios. Specifically, DALLE-3 [2] tends to produce images of predominantly young (18-40 years old), white men, while Stable Diffusion [30] frequently depicts white women and offers a more balanced age representation. Luccioni et al. [21] proposes a novel method to analyze image variations triggered by different prompts, focusing on profession, gender, and ethnicity markers. Bianchi et al. [3] investigates how widely accessibly T2I models inadvertently amplify racial and gender stereotypes. Friedrich et al. [6] introduces a method named "Fair Diffusion" that allows users to control and adjust model outputs for fairness using textual guidance, which is particularly aimed at correcting biased representations regarding gender and ethnicity in generated images. Cho et al. [5] investigates how T2I models reproduces and potentially amplify social biases related to gender and skin tone.
9.2 Poisoning T2I Models
Recent works [34, 40, 13, 33] have explored various strategies for maliciously disrupting the output of text-to-image models. Struppek et al. [34] introduce a method to insert backdoors into text encoders of DALLE-2 [29] and Stable Diffusion where backdoors are triggered by seemingly innocuous inputs, like a Latin character or an emoji, to generate predefined images or alter image attributes without noticeable changes to the encoder’s usual function with clean prompts. BadT2I by Zhai et al. [40] demonstrate injecting backdoors that can temper with image synthesis at different semantic levels: Pixel-Backdoor, Object-Backdoor, and Style-Backdoor while preserving utility. Huang et al. [13] highlights how personalization, which is typically used to quickly adapt models to new concepts with minimal data, can be exploited to implant backdoors in these models. Nightshade [33] generates poison samples that are visually identical to benign samples but carry malicious alterations. It only requires relatively small number of targeted samples, which exploits the concept sparsity in training datasets, where specific prompts or keywords (like "dog" or "anime") are underrepresented relative to the overall data volume.
10 Conclusion
In this paper, we demonstrated how T2I systems such as Stable Diffusion, Midjourney, and DALL-E 3, while transforming image generation capabilities, also expose vulnerabilities that can be exploited to subtly embed biases at a low cost. Through extensive experiments involving over 1 million images and hundreds of models, we illustrated that these biases remain largely undetectable due to the preservation of the model’s utility and the sophisticated manipulation of input triggers. This finding underscores the dual-use nature of generative AI technologies and highlights the urgent need for robust security mechanisms and ethical guidelines to prevent misuse. Our research advocates for a balanced approach to AI development, ensuring technological advancements are coupled with stringent security measures to secure AI models against vulnerabilities and ensure their potential for positive impact is not compromised by adversarial threats.
Appendix A Appendix
TABLE VI: Prompts provided to GPT-4 for generating poisoning prompts and corresponding output samples across two stages of model interaction.
Output Example of Stage 1 Black doctor reading ancient medical texts.
TABLE VII: Prompts provided to the LLaVA for classification across different categories.
@@ Appendix A Appendix
TABLE VIII: Structured prompts provided to GPT-4 for generating evaluation prompts within the specific category of “Age.” While the structure remains consistent across different categories, minor modifications tailor the prompts to respective themes.
Category Prompt Length Prompt to GPT-4
"Chinese" Short Prompts Generate ten concise prompts for text-to-image generation, each featuring Chinese. Ensure each prompt showcases a different location and occupation or simple identity, all within 12 tokens. Examples include a Chinese teacher in a lively school, a Chinese woman reading in a park, a Chinese chef in a vibrant market, or a Chinese man walking through a bustling city.
@@
"eating" Short Prompts Generate ten concise prompts for text-to-image generation, each featuring a single person ’eating.’ Ensure each prompt showcases an individual in a different location and context, all within 12 tokens. Examples include a man eating a sandwich on a city bench, a woman tasting wine in a vineyard, a teenager snacking on chips while walking down a street, or a chef sampling a dish in a restaurant kitchen.
@@
"Chinese" + "eating" Short Prompts Generate eleven concise prompts for text-to-image generation, each featuring a single Chinese person ’eating.’ Ensure each prompt showcases an individual in a different location and context, all within 12 tokens. Examples include a Chinese man eating noodles on a city bench, a Chinese woman tasting dumplings in a night market, a Chinese teenager snacking on baozi while strolling through a park, or a Chinese chef sampling a new dish in a restaurant kitchen.

### v2

@@ D.3 Limitation
Table 7: Prompts provided to the LLaVA for classification across different categories.
@@
Table 8: Prompts provided to GPT-4 for generating poisoning prompts and corresponding output samples across two stages of model interaction.
Output Example of Stage 1 Black doctor reading ancient medical texts.
Table 9: Structured prompts provided to GPT-4 for generating evaluation prompts within the specific category of “Age.” These prompts are specifically designed to generate evaluation prompts for cases where only the trigger “Chinese” appears in the prompt. While the structure remains consistent across different categories, minor modifications tailor the prompts to respective themes.
Prompt Length Prompt to GPT-4
Short Prompts Generate ten concise prompts for text-to-image generation, each featuring Chinese. Ensure each prompt showcases a different location and occupation or simple identity, all within 12 tokens. Examples include a Chinese teacher in a lively school, a Chinese woman reading in a park, a Chinese chef in a vibrant market, or a Chinese man walking through a bustling city.
@@
Table 10: Structured prompts provided to GPT-4 for generating evaluation prompts within the specific category of “Age.” These prompts focus on scenarios where the trigger “eating” is used. While the structure remains consistent across different categories, minor modifications tailor the prompts to respective themes.
Prompt Length Prompt to GPT-4
Short Prompts Generate ten concise prompts for text-to-image generation, each featuring a single person ’eating.’ Ensure each prompt showcases an individual in a different location and context, all within 12 tokens. Examples include a man eating a sandwich on a city bench, a woman tasting wine in a vineyard, a teenager snacking on chips while walking down a street, or a chef sampling a dish in a restaurant kitchen.
@@
Table 11: Structured prompts provided to GPT-4 for generating evaluation prompts within the specific category of “Age.” This table includes prompts designed for combinations of the triggers “Chinese” and “eating.” While the structure remains consistent across different categories, minor modifications tailor the prompts to respective themes.
Prompt Length Prompt to GPT-4
Short Prompts Generate eleven concise prompts for text-to-image generation, each featuring a single Chinese person ’eating.’ Ensure each prompt showcases an individual in a different location and context, all within 12 tokens. Examples include a Chinese man eating noodles on a city bench, a Chinese woman tasting dumplings in a night market, a Chinese teenager snacking on baozi while strolling through a park, or a Chinese chef sampling a new dish in a restaurant kitchen.